Reports 1-1 of 1 Clear search Modify search
MIF (General)
masayuki.nakano - 5:29 Wednesday 25 March 2020 (13799) Print this report
New recorde (around 900 kpc)

[Yokozawa, Nakano]

Thanks for the night workers, the IFO has been locked stably and got new recorde. (above 900kpc at the maximum.)
Today, the dewhitening filter of the ETMs could be engaged.

Images attached to this report
Comments to this report:
shinji.miyoki - 5:33 Wednesday 25 March 2020 (13800) Print this report
Truely Yattane! By the way, the DARM control UGF is around 100hz. This used to be 10 Hz ? Sorry for if my miss understanding about omc loop gain. If the intensity noise projection is true, we hope more but slight improvement of ISS contributes.
takaaki.yokozawa - 6:48 Wednesday 25 March 2020 (13803) Print this report

Quick noise investigation by Bruco.

1. MICH and PRCL control noise is smaller than before. If alignment is good, the effect from them became smaller.(Fig.1. and Fig.2.)

2. 206Hz. As you can see, there is coherence with CARM and some IMC channels(Fig.3.). This is just assumption, the IMC alignment may affect this peak. (Amplithde is varied everyday)

3. There is coherence with peaks near at 120Hz and 240Hz in the IMTX's and ITMY's PS signals and magnetometers. We may investigate the power supply around them.

4. I couldn't find the clear reason for 90Hz peak

Images attached to this comment
tomotada.akutsu - 8:35 Wednesday 25 March 2020 (13804) Print this report

Bravo! Oh... Bravo!!

Search Help